24/10+ m +1=24/10-m
What would be the simplest way to solve this equation?

Answer by ewatrrr(24785)

 
Hi,
24/10+ m +1=24/10-m
What would be the simplest way to solve this equation?
combine Like terms..Isolate the variable on one side of the EQ
24/10+ m +1=24/10-m ||Adding m and -24/10 and -1 to both sides of the EQ
2m = -1
m = -1/2